# Rigor in AI: Doing Rigorous AI Work Requires a Broader, Responsible AI-Informed Conception of Rigor

### v1

@@ 2.5 Reporting rigor
The understanding of research findings depends on what is communicated about these findings and how. Reporting rigor is concerned with making sure research findings are clearly and appropriately communicated and justified. For instance, assume a researcher wishes to compare the performance of different recommender systems. Even when reporting only aggregated results, multiple options are possible, including averaging across ratings (treating each rating as equally important regardless of which user provided it or what item it was provided for), users (treating each user equally by first computing performance at user level), or items (treating each item equally by first computing performance at item level). [^3] While this may often be left implicit due to a failure to explicitly systematize what performance entails (§2.3), each option likely reflects a somewhat different underlying conceptualization of performance—e.g., average user, item, rating-level performance—with “[d]ifferent ‘findings’ created by different conceptualizations” [54]. Depending on the data distribution (e.g., number of ratings per item/user), these differences may lead someone to draw different or even contradictory conclusions about which model performs best [e.g., 105]. Some ratings may also be more difficult to predict than others [e.g., 117], and any aggregation can obfuscate where exactly the model fails or succeeds [31, 117, 87]. Further, even such simple aggregations can introduce tacit assumptions about what should be optimized for e.g., to ensure a good predictive performance across all users versus all items [31]. Making these choices explicit can facilitate others’ understanding of what specifically is being reported and why.
Such choices of what findings to report, and how—much as with choices of research questions, theoretical constructs, or methods—are shaped by our beliefs, values, and preferences, as well as disciplinary norms and incentives. Negative results are, for instance, less likely to be reported or published [133, 150], potentially “lead[ing] others to develop overly optimistic ideas about scientific progress on a particular topic” [10], and what is reported may be cherry-picked, such as “uncommonly compelling examples to illustrate the output of generative models” [10]. And as the example above also illustrates, even for the same findings, how the findings are reported or what about the findings is reported matters—such as choosing to report inferential uncertainty (“how precisely we have estimated the average for each group,” when interested in estimating aggregate outcomes) versus outcome variability (“how much individual outcomes vary around averages for each group”) [154]. Communicating the former risks leading people to “overestimate the size and importance of scientific findings” [154].
@@
Mechanisms to promote reporting rigor Pre-study practices around documenting and disclosing how a study would be run and what about it would be reported like pre-registration [e.g., 64, 103, 137], and around reporting more granular, disaggregated results [e.g., 14, 31] can promote reporting rigor.
Pre-registration and reporting practices:  Reflecting on what to report about a study before the study is even conducted can help mitigate concerns about making such choices post-factum by hypothesizing after the results are known or selectively presenting only from positive results [76, 119, 103]. Pre-registration is the “practice of specifying what you are going to do, and what you expect to find in your study, before carrying out the study” [137]. While pre-registration is seen as a mechanism for promoting more reliable and replicable research findings by differentiating between confirmatory research—where hypotheses are tested and pre-registration is required—and exploratory research—where hypotheses are generated and pre-registration is not required [132, 103], it also asks us to reflect on and pre-set what metrics and measures of success will be used and reported on to minimize post-factum hypothesizing [64].
Disaggregated evaluations:  As illustrated by the earlier example, aggregate measurements and metrics can obscure rare phenomena and information about where systems or models tend to fail or succeed, and can mask important effects [129, 31, 14, 117]. To mitigate such concerns, many have called for reporting disaggregated evaluations [31, 87, 68], which “have proven to be remarkably effective at uncovering the ways in which AI systems perform differently for different groups of people” [14] and deemed a “critical piece of full empirical analysis” [129]. While conceptually simple, “their results, conclusions, and impacts depend on a variety of choices” [14], and they require careful consideration and justification—including by engaging with domain experts [135]—of how different choices of why, when, what, and how to conduct and report on such evaluations shape what inferences can be drawn and their impact. This however can in turn help make such choices explicit and encourage debates about their appropriateness.
@@ 2.6 Interpretative rigor
Assume an AI system is used to solve International Math Olympiad (IMO) problems due to achieving high accuracy on a benchmark designed to assess mathematical reasoning [e.g., 120, 55]. As Salaudeen et al. [120] note, based on the system’s performance on the benchmark, two different alternative claims could be considered: the system can “solve linear algebra questions from a textbook accurately” or the system “has reached human-expert-level mathematical reasoning.” Reliably moving from performance on a benchmark to any of the two claims requires clarity about any background assumptions concerning the feasibility of an AI system reaching human-level reasoning abilities [e.g., 138] (epistemic and normative rigor, §2.12.2), about how both “mathematical reasoning” and “human-expert-level” are conceptualized (conceptual rigor, §2.3), about whether the benchmark actually measures mathematical reasoning ability (methodological rigor, §2.4), and about how findings were reported—e.g., do we know what the performance on linear algebra questions is? (reporting rigor, §2.5). As others have noted, a pernicious trap “is to believe that [methods] bestow a natural interpretative clarity and self-reflexive awareness on the researcher” [95], and thus “scientists must acknowledge the social and interpretative character of scientific discovery” [22].
@@
Both making and understanding knowledge claims require interpretation. There are often multiple perspectives through which empirical, experimental, or theoretical evidence could be interpreted, and these different perspectives may not only lead to different aspects being emphasized but also lead to different or even contradictory conclusions being drawn. While reporting rigor2.5) is concerned with choices about what findings to report and how, interpretative rigor is concerned with the choices we make when we move from findings to descriptive claims—e.g., the system accurately solves a given task—or to prescriptive or normative claims—e.g., the system should be used to replace humans. Such claims rarely just directly follow from findings, but are shaped by choices and considerations related to all other facets of rigor (§2.12.4)—they are influenced and made in the context of background knowledge, the relationship with the theoretical constructs under use, and the methods used to produce the findings and their limitations. While critical to how any work ultimately intervenes in the world, how claims are arrived at is often overlooked in discussions about rigor in AI, with the interpretation of results—i.e., what they mean, what people should do next—being treated as self-evident. The criteria for interpretative rigor is also “not whether the same interpretation would be independently arrived upon by different” people, but rather that “based on the evidence provided, is a given interpretation credible” or if “given all the same source information, would the interpretation stand up to scrutiny as being a justified, empirically grounded, exposition of the phenomenon” [104].
@@
Mechanisms to promote interpretative rigor Documenting and justifying evidence informing or situating possible claims can promote interpretative rigor, like via documenting AI-related artifacts [e.g., 15, 100, 52] or by engaging with possible threats to internal/external validity [e.g., 131, 85, 106].
@@
Documenting AI-related artifacts, their limitations, and their impacts:  Transparency about any AI-related artifacts (e.g., datasets, models, systems) under use can facilitate others’ understanding of what claims may or may not be possible by providing added context about their characteristics and intended uses. To help scaffold and promote more transparent reporting on AI-related artifacts, researchers have developed tools, resources, or what Boyd [28] terms “context documents” (e.g., for datasets [111, 15, 52], models [100], services [113]). Complementing these efforts, others have argued for and put forward practices for recognizing and disclosing the limitations—i.e., “drawbacks in the design or execution of research that may impact the resulting findings and claims” [131]—and impacts—i.e., actual or possible consequences from the research, development, deployment, or use—of AI artifacts [27, 131, 94, 72, 11, 46, 88, 107, 98]. Limitations, in particular, directly impact how research findings can be interpreted—a failure to recognize them can lead to unsubstantiated claims, while a failure to disclose them can further lead to misinterpretations or misuse of claims. Impacts, in turn, can further affect prescriptive and normative claims and their implications, such about how one should act given the research findings.
@@
Internal and external validity:  Interpretative rigor requires not only careful deliberation on what the claims are about, but also whether there is appropriate evidence to support the claims. Establishing whether the research findings constitute appropriate evidence requires engaging with both internal validity—whether there are unaddressed issues with study design or execution that may compromise the findings such as data leakage [e.g., 75] or improper baseline comparisons [e.g., 85]—and external validity—whether the findings generalize to different settings such as from one dataset [e.g., 110] or construct [e.g., 144] to another. For broader discussions, see Olteanu et al. [106] and Liao et al. [85].
@@
By making the case for better documentation [100, 113, 52], evaluation practices [143, 155, 68], development and deployment practices [12, 72, 112], and understanding of impacts and limitations [27, 131, 11], responsible AI research asks for greater scientific rigor. The AI community has too often cast responsible AI considerations as out of scope, but holds up research rigor as a virtue. In AI research and practice, however, rigor remains largely understood in terms of methodological rigor. Nevertheless, this can have unintended consequences as, for instance, “if the methodology is considered to be the sine qua non of scientificity, as it usually is, then there will be enormous pressures for the structure of all theories to accommodate to the theoretical structure embedded in the methodology [… with each] embedded theory involv[ing] its own value hierarchy” (emphasis original) [42].
@@
We argue that rigor in AI means more than just methodological rigor, and in so doing we bring responsible AI as it pertains to six distinct aspects of rigor—epistemic, normative, conceptual, methodological, reporting, and interpretative—under the umbrella of an AI researcher and practitioner’s responsibility. By calling attention to these different facets of rigor, we also hope to provide the AI community with useful language that can help it raise, clarify, and examine a wider range of concerns about existing practices in AI work. Nevertheless, while a broader conception of rigor can improve research integrity and quality, rigor is not a panacea for all problems in AI research and practice. We do not claim that these facets are all-inclusive, but that they help demonstrate how expanding our conception of rigor beyond methodological considerations can help contribute evidence that work is rigorous.
Alternative Views
@@
AI work is already rigorous or rigorous enough:  Some may view addressing methodological rigor concerns as sufficient for ensuring rigorous AI work; under this view, rigor equates to and is achieved by focusing on methodological concerns. As underscored throughout this paper, this view leaves unaddressed a range of concerns which have produced undesirable outcomes, including a reliance on pseudo-scientific assumptions [e.g., 134], treatment of social phenomena inconsistent with broader scholarly understanding [e.g., gender 44, 77], systems that are not fit-for-purpose or cause harm [e.g., 112, 69], and claims or use of language that impedes public understanding of AI [e.g., 37].
@@
Non-methodological concerns are outside the purview of AI work:  This view may arise because AI researchers and practitioners may see such concerns as outside the scope of core AI work [155]; they may also not see themselves as well-suited to addressing such concerns, either because it would be difficult to acquire the expertise needed, or because some concerns ought instead to be addressed by subject matter experts (with whom they may not have the time, desire, or resources to engage). We agree that engaging with subject matter experts can be valuable, and believe that AI work has been strengthened when it has done so [143]. Nevertheless, since all work involves choices about what problems are important, why they are important, and what precise objects are under investigation, and since AI work often claims real-world impact in or relevance to particular domains, we argue that it is impossible to do any AI work that does not implicate epistemic, normative, or conceptual questions—and thus researchers and practitioners must grapple with these concerns explicitly rather than implicitly.
@@
All rigor concerns are validity concerns:  Under this view, this presentation of rigor concerns simply reframes work already addressing well-understood validity threats in AI research [e.g., 106, 71, 143, 85, 120]. While some of the concerns and mechanisms we describe (e.g., construct clarity/validity, internal/external validity) appear in the literature on validity, many considerations, particularly those related to epistemic, normative, conceptual, and interpretative rigor, precede or are a prerequisite to questioning and establishing validity, and facilitate reflection about issues beyond validity. For example, a failure to interrogate one’s epistemological and normative commitments may result in systems that operationalize a construct as defined, but whose definition has been contested.
@@ Acknowledgments
We thank Micheal Veale and Hanna Wallach for early conversations that have motivated this paper. We are also grateful to the members of the STAC team at Microsoft Research NYC for their feedback.

### v2

@@ 2.5 Reporting rigor
The understanding of research findings depends on what is communicated about these findings and how. Reporting rigor is concerned with making sure research findings are clearly and appropriately communicated and justified. For instance, assume a researcher wishes to compare the performance of different recommender systems. Even when reporting only aggregated results, multiple options are possible, including averaging across ratings (treating each rating as equally important regardless of which user provided it or what item it was provided for), users (treating each user equally by first computing performance at user level), or items (treating each item equally by first computing performance at item level). [^4] While this may often be left implicit due to a failure to explicitly systematize what performance entails (§2.3), each option likely reflects a somewhat different underlying conceptualization of performance—e.g., average user, item, rating-level performance—with “[d]ifferent ‘findings’ created by different conceptualizations” [george1973conceptualization]. Depending on the data distribution (e.g., number of ratings per item/user), these differences may lead someone to draw different or even contradictory conclusions about which model performs best [e.g., olteanu2014comparing]. Some ratings may also be more difficult to predict than others [e.g., rodriguez2021evaluation], and any aggregation can obfuscate where exactly the model fails or succeeds [burnell2023rethink, rodriguez2021evaluation, lipton2019troubling]. Further, even such simple aggregations can introduce tacit assumptions about what should be optimized for e.g., to ensure a good predictive performance across all users versus all items [burnell2023rethink]. Making these choices explicit can facilitate others’ understanding of what specifically is being reported and why.
Such choices of what findings to report, and how—much as with choices of research questions, theoretical constructs, or methods—are shaped by our beliefs, values, and preferences, as well as disciplinary norms and incentives. Negative results are, for instance, less likely to be reported or published [song2010dissemination, wilholt2009bias], potentially “lead[ing] others to develop overly optimistic ideas about scientific progress on a particular topic” [ashurst2022disentangling], and what is reported may be cherry-picked, such as “uncommonly compelling examples to illustrate the output of generative models” [ashurst2022disentangling]. And as the example above also illustrates, even for the same findings, how the findings are reported or what about the findings is reported matters—such as choosing to report inferential uncertainty (“how precisely we have estimated the average for each group,” when interested in estimating aggregate outcomes) versus outcome variability (“how much individual outcomes vary around averages for each group”) [zhang2023illusion]. Communicating the former risks leading people to “overestimate the size and importance of scientific findings” [zhang2023illusion].
@@
Mechanisms to promote reporting rigor Pre-study practices around disclosing how a study would be run and what about it would be reported like pre-registration [e.g., hofman2023pre, nosek2018preregistration, van2021preregistering], and around reporting more granular, disaggregated results [e.g., mitchell2019model, barocas2021designing, burnell2023rethink] can promote reporting rigor.
Pre-registration and reporting practices:  Reflecting on what to report about a study before conducting it can mitigate concerns about making such choices post-factum by hypothesizing after the results are known or reporting only from positive results [kerr1998harking, rubin2017does, nosek2018preregistration]. Pre-registration is considered the “practice of specifying what you are going to do, and what you expect to find in your study, before carrying out the study” [van2021preregistering]. Pre-registration is seen as a mechanism for promoting more reliable research findings by differentiating between confirmatory—where hypotheses are tested and pre-registration is required—and exploratory research—where hypotheses are generated and pre-registration may not be required [sogaard2023two, nosek2018preregistration]. While pre-registration often relies on narrow notions of what constitutes exploratory research and might inappropriately situate confirmatory research as more reliable [feest2025toward], as a mechanism it can be used not only to encourage reflection on what measures of success will be used and reported on but also on how the study is designed [hofman2023pre], including in exploratory settings.
Disaggregated evaluations:  As illustrated by the earlier example, aggregate measurements and metrics can obscure rare phenomena and information about where systems or models tend to fail or succeed, and can mask important effects [sculley2018winner, burnell2023rethink, barocas2021designing, rodriguez2021evaluation]. To mitigate such concerns, many have called for reporting disaggregated evaluations [mitchell2019model, burnell2023rethink, lipton2019troubling, hutchinson2022evaluation], which “have proven to be remarkably effective at uncovering the ways in which AI systems perform differently for different groups of people” [barocas2021designing] and deemed a “critical piece of full empirical analysis” [sculley2018winner]. While conceptually simple, “their results, conclusions, and impacts depend on a variety of choices” [barocas2021designing], and they require careful consideration and justification—including by engaging with domain experts [stevens2020recommendations]—of how different choices of why, when, what, and how to conduct and report on such evaluations shape what inferences can be drawn and their impact. This, however, can in turn help make such choices explicit and encourage debates about their appropriateness.
@@ 2.6 Interpretative rigor
Assume an AI system achieves high accuracy on a benchmark designed to assess mathematical reasoning [e.g., salaudeen2025measurement, glazer2024frontiermath]. As salaudeen2025measurement note, based on the system’s performance on the benchmark, both of these two different alternative claims could be considered: the system can “solve linear algebra questions from a textbook accurately” or the system “has reached human-expert-level mathematical reasoning.” Reliably moving from performance on a benchmark to either of the two claims requires clarity about any background assumptions concerning the feasibility of an AI system reaching human-level reasoning abilities [e.g., van2024reclaiming] (epistemic and normative rigor, §2.12.2), about how both “mathematical reasoning” and “human-expert-level” are conceptualized (conceptual rigor, §2.3), about whether the benchmark actually measures mathematical reasoning ability (methodological rigor, §2.4), and about how findings were reported—e.g., do we know what the performance on linear algebra questions is? (reporting rigor, §2.5). Echoing markham2009response’s observation, a pernicious trap is to believe that methods and evaluation practices “bestow a natural interpretive clarity and self-reflexive awareness on the researcher,” and thus “scientists must acknowledge the social and interpretative character of scientific discovery” [birhane2023science].
@@
Both making and understanding knowledge claims require interpretation. There are often multiple perspectives through which empirical, experimental, or theoretical evidence could be interpreted, and these different perspectives may not only lead to different aspects being emphasized but may also lead to different or even contradictory conclusions being drawn. While reporting rigor2.5) is concerned with choices about what findings to report and how, interpretative rigor is concerned with the conclusions we draw from these findings, and thus with the choices we make when we move from findings to either some descriptive claims—e.g., the system solves a given task—or to some prescriptive or normative claims—e.g., the system should be used to replace humans. Such conclusions or claims rarely directly follow from findings, but require some interpretation and are shaped by choices and considerations related to all other facets of rigor (§2.12.4)—they are influenced and made in the context of background knowledge, the relationship with the theoretical constructs under use, and the methods used to produce the findings and their limitations. While critical to how any work ultimately intervenes in the world, how claims are arrived at is often overlooked in discussions about rigor in AI, with the interpretation of results—i.e., what they mean, what people should do next—being treated as self-evident. The criteria for interpretative rigor is also “not whether the same interpretation would be independently arrived upon by different” people, but rather that “based on the evidence provided, is a given interpretation credible” or if “given all the same source information, would the interpretation stand up to scrutiny as being a justified, empirically grounded, exposition of the phenomenon” [nowell2019reviewer].
@@
Mechanisms to promote interpretative rigor Documenting and justifying evidence informing or situating possible claims can promote interpretative rigor, like via documenting AI artifacts [e.g., bender2018data, mitchell2019model, gebru2021datasheets] or by engaging with possible threats to internal/external validity [e.g., smith2022real, liao2021we, olteanu2019social].
@@
Documenting AI artifacts, their limitations, and their impacts:  Transparency about any AI artifacts (e.g., datasets, models, systems) under use can facilitate others’ understanding of what claims may or may not be possible by providing added context about their characteristics and intended uses. To help scaffold and promote more transparent reporting on AI artifacts, researchers have developed tools, resources, or what boyd2021datasheets terms “context documents” (e.g., for datasets [pushkarna2022data, bender2018data, gebru2021datasheets], models [mitchell2019model], services [arnold2019factsheets]). Complementing these efforts, others have argued for and put forward practices for recognizing and disclosing the limitations—i.e., “drawbacks in the design or execution of research that may impact the resulting findings and claims” [smith2022real]—and impacts—i.e., actual or possible consequences from the research, development, deployment, or use—of AI artifacts [boyarskaya2020overcoming, smith2022real, malik2020hierarchy, javadi2021monitoring, ashurst2022ai, do2023s, liu2022examining, olteanu2023responsible, metcalf2021algorithmic]. Limitations, in particular, directly impact how research findings can be interpreted—a failure to recognize them can lead to unsubstantiated claims, while a failure to disclose them can further lead to misinterpretations or misuse of claims. Impacts, in turn, can further affect prescriptive and normative claims and their implications, such about how one should act given the research findings.
@@
Internal and external validity:  Interpretative rigor requires not only careful deliberation on what the claims are about, but also whether there is appropriate evidence to support the claims. Establishing whether the research findings constitute appropriate evidence requires engaging with both internal validity—whether there are unaddressed issues with study design or execution that may compromise the findings such as data leakage [e.g., kaufman2012leakage] or improper baseline comparisons [e.g., liao2021we]—and external validity—whether the findings generalize to different settings such as from one dataset [e.g., porada2024challenges] or construct [e.g., wang2025identities] to another. For broader discussions, see olteanu2019social and liao2021we.
@@
By making the case for better documentation [mitchell2019model, arnold2019factsheets, gebru2021datasheets], better evaluation practices [wallach2025position, zhou2022deconstructing, hutchinson2022evaluation], better development and deployment practices [balayn2023faulty, javadi2021monitoring, raji2022fallacy], and a better understanding of impacts and limitations [boyarskaya2020overcoming, smith2022real, ashurst2022ai], responsible AI research asks for greater scientific rigor. The AI community has too often cast responsible AI considerations as out of scope, but holds up research rigor as a virtue. In AI research and practice, however, rigor still remains largely understood in terms of methodological rigor. Nevertheless, this can have unintended consequences as, for instance, “if the methodology is considered to be the sine qua non of scientificity, as it usually is, then there will be enormous pressures for the structure of all theories to accommodate to the theoretical structure embedded in the methodology [… with each] embedded theory involv[ing] its own value hierarchy” (emphasis original) [danziger1985methodological]. In our reconception of rigor in AI—which expands beyond merely methodological rigor—we reframe many calls of the responsible AI community as within scope of all AI researchers and practitioners.
@@
We argue that rigor in AI research and practice means more than just methodological rigor, and in so doing we bring doing AI work responsibly—as it pertains to epistemic, normative, conceptual, methodological, reporting, and interpretative rigor—under the umbrella of an AI researcher and practitioner’s responsibility. By calling attention to these different facets of rigor, we also hope to provide the AI community with useful language that can help researchers and practitioners raise, clarify, and examine a wider range of concerns about existing practices in AI work. Nevertheless, while a broader conception of rigor can improve research integrity and quality, rigor is not a panacea for all problems in AI research and practice. While the facets of rigor we foreground are foundational to good science and apply broadly, we grounded our discussion about these facets in critical work from the responsible AI community because that work discusses these aspects in the context of AI research and practice—which is what our call is concerned with. We also do not claim that these facets are all-inclusive, but rather that they help demonstrate how expanding our conception of rigor beyond methodological considerations can help contribute evidence that AI work is rigorous.
Alternative Views on Rigor in AI Research and Practice
@@
AI work is already rigorous or rigorous enough:  Some may view addressing methodological rigor concerns as sufficient for ensuring rigorous AI work; under this view, rigor equates to and is achieved by focusing on methodological concerns. As underscored throughout this paper, this view leaves unaddressed a range of concerns which have produced undesirable outcomes, including a reliance on pseudo-scientific assumptions [e.g., stark2021physiognomic], treatment of social phenomena inconsistent with broader scholarly understanding [e.g., gender devinney2022theories, keyes2018misgendering], systems that are not fit-for-purpose or cause harm [e.g., raji2022fallacy, hutiri2024not], and claims or use of language that impedes public understanding of AI [e.g., cheng2024one].
@@
Non-methodological concerns are outside the purview of AI work:  This view may arise because AI researchers and practitioners may see such concerns as outside the scope of core AI work [zhou2022deconstructing]; they may also not see themselves as well-suited to addressing such concerns, either because it would be difficult to acquire the expertise needed, or because some concerns ought instead to be addressed by subject matter experts (with whom they may not have the time, desire, or resources to engage). We agree that engaging with subject matter experts can be valuable, and believe that AI work has been strengthened when it has done so [wallach2025position]. Nevertheless, since all work involves choices about what problems are important, why they are important, and what precise objects are under investigation, and since AI work often claims real-world impact in or relevance to particular domains, we argue that it is impossible to do any AI work that does not implicate epistemic, normative, or conceptual questions—and thus researchers and practitioners must grapple with these concerns explicitly rather than implicitly.
@@
All rigor concerns are validity concerns:  Under this view, this presentation of rigor concerns simply reframes work already addressing well-understood validity threats in AI research [e.g., olteanu2019social, jacobs2021measurement, wallach2025position, liao2021we, salaudeen2025measurement]. While some of the concerns and mechanisms we describe (e.g., construct clarity/validity, internal/external validity) appear in the literature on validity, many considerations, particularly those related to epistemic, normative, conceptual, and interpretative rigor, precede or are a prerequisite to questioning and establishing validity, and facilitate reflection about issues beyond validity. For example, a failure to interrogate one’s epistemological and normative commitments may result in systems that operationalize a construct as defined, but whose definition has been contested.
@@ Acknowledgments
We thank Michael Veale and Hanna Wallach for early conversations that have motivated this paper. We are also grateful to the members of the STAC team at Microsoft Research NYC for their feedback.